# Detectability of dark matter subhalo impacts in Milky Way stellar streams

### v2

@@ 4.3 Confidence intervals on subhalo mass
In the denominator, $\hat{\bm{\theta}}$ refers to the values of the nuisance parameters which maximize the likelihood for that particular value of $M_{\rm sh}$. By definition, $t=0$ at the maximum likelihood point. In the large sample limit, the maximum likelihood point will be at $M_{\mathrm{sh}}^{\mathrm{true}}$ and the $n\sigma$ confidence interval on $M_{\mathrm{sh}}^{\mathrm{true}}$ is obtained by solving for the $M_{\rm sh}$ at which $t=n^{2}$.
@@
Figure 11: Test statistic $-t$ as a function of $M_{\mathrm{sh}}$ used to obtain confidence intervals, as defined in Eq. 28. Here $-t$ is evaluated on the Asimov data set for a stream with $\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$, $\lambda=100~{}\mathrm{deg}^{-1}$ and all the four observational scenarios. The dataset is generated from $M_{\mathrm{sh}}^{\mathrm{true}}=10^{8}\mathrm{M}_{\odot}$.
@@
Fig. 11 shows $-t$ as a function of $M_{\mathrm{sh}}$, evaluated on the Asimov data set. As in the case for $q_{0}$, we evaluate $t(M_{\mathrm{sh}})$ using scipy.optimize.minimize to find the maximized likelihood $L(M_{\mathrm{sh}},~{}\hat{\bm{\theta}})$ as a function of $M_{\mathrm{sh}}$. We determine the $2\sigma$ confidence interval on $M_{\mathrm{sh}}^{\mathrm{true}}$ from the intersection of the lines with $-t=-n^{2}=-4$ (dashed line). We see that the confidence interval spans further on the lower mass end compared to the higher mass end. This is expected because the subhalo is more difficult to detect at the low subhalo mass end, leading to higher uncertainties and wider confidence intervals. The opposite is true for the high subhalo mass end.
@@ 4.4 Minimal stream length in angle
In addition to the stream properties used in Eq. 27, the stream length in angle, denoted by $l$, also affects detectability. So far, the default angular region in the likelihood calculations is $[-22^{\circ},22^{\circ}]$ around the impact location, for a total range of $\Delta\psi=44^{\circ}$ on a single stream arm. This region was selected to be able to capture the full perturbation profile of a fairly massive subhalo as high as $\sim 10^{9}~{}\mathrm{M}_{\odot}$. However, this assumes a stream which is at least twice the angular length, which would impose a strong requirement when considering Milky Way streams.
On the other hand, for the minimum detectable subhalo mass $M_{\rm sh}^{\rm min}$, the requirement on the stream length can be reduced significantly. We now consider variable angular regions of size $[-\Delta\psi/2,\Delta\psi/2]$ with total range $\Delta\psi$. In Fig. 13, we plot the median $M_{\rm sh}^{\rm min}$ as a function of $\Delta\psi$ for varying stream properties $\sigma_{\theta}$, $r_{0}$, and $\lambda$. For simplicity, we only consider the best observational scenario “Via + Gaia + LSST10” here. We follow the same procedure as that used to obtain Fig. 8. We only show a data point if the resulting $M_{\rm sh}^{\rm min}$ is smaller than $10^{8}~{}\mathrm{M}_{\odot}$.
@@ 4.5 Dependence on impact properties
  • Impact parameter $b$ Our default value of $b=0$ is the optimal case. The results for non-zero $b$ are shown in the top left panel of Fig. 14. The detectability remains almost the same when the impact parameter $b$ is below 100 pc, and then begins to degrade when $b$ goes beyond 100 pc. As a reference, the example stream we use here has a 1-sigma width of 35 pc. Physically, the effect of the impact will remain relatively constant as long as $b$ is smaller than or comparable to the stream width. For thicker streams, the minimum subhalo mass should remain constant for even larger $b$. Our results are limited to $b\leq 10^{3}$ pc given the assumptions of the analytic approximation, Eq. 4.
@@
    Figure 15: Decomposition of $q_{0}$ contributions from different observables. All settings are the same as the top panel of Fig. 9 except that we consider special times since flyby, $t=279.3$ Myr (top) and $t=558.6$ Myr (bottom), when the perturbation in the $z$ direction is 0. The contribution from $\overline{\Delta\theta}$ vanishes completely in these cases. Subhalo masses are chosen to be the minimum detetctable subhalo mass, and $q_{0}$ is evaluated on the Asimov data set.
    As $t$ increases, the variations with $t$ become smaller. This is because the size of the gap caused by a subhalo impact grows with time, which makes the density observable $\rho/\rho_{0}$ more important over other periodic observables for detectability. The gap growth also causes a slight decrease in $M_{\rm sh}^{\rm min}$ as larger gaps are easier to detect. Fig. 15 also shows how the density observable plays a more important role as $t$ grows.
  • Direction of subhalo velocity $\hat{w}$ In studying the effects of changing the absolute subhalo velocity $\bm{w}=(w_{r},w_{t},w_{z})$, we first keep its magnitude fixed at $w=180$ km/s, and look into the effect of different directions $\hat{w}$. Taking symmetry considerations into account, we pick the following subhalo directions as representatives: $\hat{z}~{}\mathrm{(default)},\hat{r},\pm\hat{t}+0.1\hat{z},\hat{z}+\hat{r},\hat{z}\pm\hat{t}$ and $\hat{r}\pm\hat{t}$, where $\hat{t}$ is direction along the stream motion. Note that velocities purely along ±$\pm\hat{t}$ break the assumption of the analytic model, Eq. 4, so we have added a small perpendicular component $0.1\hat{z}$ to them. The results are shown in the top right panel of Fig. 14.
@@ 5 Application to known streams
Name $\sigma_{\theta}$ $l$ $r_{h}$ $M_{\mathrm{stellar}}$ $\lambda$ Retro/ $M_{\mathrm{sh}}^{\mathrm{min}}$ [$\mathrm{M}_{\odot}$] $M_{\mathrm{sh}}^{\mathrm{min}}$ [$\mathrm{M}_{\odot}$] $M_{\mathrm{sh}}^{\mathrm{min}}$ [$\mathrm{M}_{\odot}$] $M_{\mathrm{sh}}^{\mathrm{min}}$ [$\mathrm{M}_{\odot}$]
[] [] [kpc] [$\mathrm{M}_{\odot}$] [deg-1] Prograde Gaia DESI + Gaia Via + LSST Via + LSST10
C-12 0.26 28 11.5 14000 534 P 3.01e+06 2.99e+06 6.73e+05 3.97e+05
ATLAS-Aliqa Uma 0.13 41 21.4 19000 495 P 4.59e+06 4.65e+06 7.76e+05 3.99e+05
300S 0.17 25 15.9 7600 325 R 5.06e+06 5.13e+06 9.90e+05 5.42e+05
NGC 6397 0.39 32 2.5 2500 84 P 1.22e+06 1.24e+06 6.40e+05 5.77e+05
GD-1 0.21 119 8.0 14000 126 R 3.83e+06 3.93e+06 1.10e+06 7.23e+05
Palomar 5 0.27 32 21.3 17000 568 P 9.68e+06 9.57e+06 1.51e+06 7.27e+05
Orphan-Chenab 0.51 137 20.7 130000 1014 P 1.21e+07 1.16e+07 1.78e+06 8.23e+05
Ylgr 0.36 49 9.5 11000 240 R 5.93e+06 5.92e+06 1.44e+06 8.63e+05
Kshir 0.11 37 10.7 2200 64 R 5.42e+06 5.73e+06 1.42e+06 8.93e+05
Gaia-6 0.2 21 8.3 1800 92 R 5.03e+06 5.20e+06 1.42e+06 9.25e+05
NGC 5466 0.11 23 17.4 1900 88 R 1.04e+07 1.10e+07 2.06e+06 1.11e+06
C-7 0.21 34 5.8 1500 47 R 4.64e+06 4.86e+06 1.61e+06 1.16e+06
Gaia-1 0.17 40 5.0 1100 29 R 4.02e+06 4.26e+06 1.57e+06 1.19e+06
Jhelum 0.33 97 13.0 17000 187 P 1.19e+07 1.20e+07 2.44e+06 1.33e+06
Leiptr 0.24 73 7.1 3000 44 R 8.49e+06 8.84e+06 2.59e+06 1.72e+06
Gaia-8 0.41 47 7.1 3600 82 R 9.61e+06 9.74e+06 2.74e+06 1.75e+06
Gaia-12 0.16 29 11.4 1200 44 R 1.29e+07 1.35e+07 3.12e+06 1.84e+06
SGP-S 0.1 29 9.5 530 20 P 1.02e+07 1.10e+07 2.92e+06 1.88e+06
New-25 0.44 23 50.3 18000 837 P 6.47e+07 6.23e+07 5.77e+06 2.02e+06
C-19 0.18 47 18.0 3200 73 P 2.34e+07 2.43e+07 4.29e+06 2.18e+06
Hrid 0.5 77 3.2 2000 28 R 6.47e+06 6.66e+06 2.89e+06 2.29e+06
Kwando 3.29 57 7.7 40000 750 P 2.16e+07 1.98e+07 4.53e+06 2.42e+06
Omega Centauri 2.13 81 5.2 20400 269 R 1.41e+07 1.34e+07 3.95e+06 2.45e+06
Sylgr 0.49 30 3.3 700 25 P 7.33e+06 7.57e+06 3.22e+06 2.54e+06
C-13 0.25 23 8.4 750 35 P 1.48e+07 1.55e+07 4.09e+06 2.54e+06
Phlegethon 0.46 77 3.4 1700 24 R 7.52e+06 7.79e+06 3.28e+06 2.56e+06
Indus 0.99 90 16.6 34000 404 P 3.73e+07 3.57e+07 5.79e+06 2.66e+06
C-24 0.29 37 14.8 2700 78 P 2.69e+07 2.75e+07 5.21e+06 2.70e+06
C-23 0.16 20 8.7 310 17 P 1.72e+07 1.84e+07 4.91e+06 3.11e+06
New-16 0.17 22 5.0 170 8 P 1.13e+07 1.23e+07 4.47e+06 3.32e+06
NGC 3201 0.38 111 4.9 2100 20 R 1.37e+07 1.43e+07 4.93e+06 3.48e+06
Slidr 0.83 35 3.1 840 26 P 1.19e+07 1.20e+07 5.08e+06 3.89e+06
New-1 0.1 31 15.5 420 14 P 3.30e+07 3.59e+07 7.14e+06 3.90e+06
NGC 288 0.62 30 11.0 2200 78 R 3.75e+07 3.75e+07 7.91e+06 4.21e+06
Gaia-9 0.75 36 4.2 950 28 R 1.74e+07 1.77e+07 6.32e+06 4.43e+06
M5 0.16 39 14.2 710 20 P 3.82e+07 4.08e+07 8.23e+06 4.45e+06
Gaia-11 0.23 50 12.5 1100 24 R 3.96e+07 4.17e+07 8.79e+06 4.81e+06
C-11 0.74 33 6.9 1100 36 P 3.64e+07 3.67e+07 9.91e+06 5.94e+06
M3 0.48 76 9.8 2000 28 P 5.15e+07 5.29e+07 1.21e+07 6.70e+06
New-7 0.38 26 2.8 110 4 P 1.61e+07 1.73e+07 8.15e+06 6.70e+06
C-9 0.86 25 5.4 630 27 P 3.43e+07 3.47e+07 1.06e+07 6.76e+06
New-10 0.5 26 2.3 100 4 P 1.66e+07 1.77e+07 9.11e+06 7.76e+06
M68 0.69 105 5.4 1700 17 P 3.79e+07 3.89e+07 1.21e+07 7.79e+06
New-13 0.38 29 6.2 250 9 R 4.08e+07 4.31e+07 1.29e+07 8.32e+06
New-23 0.34 30 4.2 120 4 R 3.18e+07 3.42e+07 1.29e+07 9.36e+06
C-25 1.06 25 11.0 1300 56 P 9.39e+07 9.30e+07 1.87e+07 9.39e+06
Wukong 3.0 63 18.1 15000 254 P 2.39e+08 2.23e+08 3.13e+07 1.25e+07
New-3 1.08 71 1.7 120 2 P 4.58e+07 4.84e+07 2.76e+07 2.37e+07
New-6 1.27 80 3.1 340 4 P 8.21e+07 8.49e+07 3.40e+07 2.42e+07
New-2 1.73 61 2.2 73 1 P 1.71e+08 1.79e+08 8.42e+07 6.34e+07
@@
Table 5: Milky Way streams and estimated minimum detectable subhalo mass at 95% CL under different observational scenarios. Stream properties are obtained from the catalog in Bonaca & Price-Whelan (2025). We only include streams with reported stellar mass, angular length $l>20^{\circ}$, and density of observable stars greater than 1 star per $2^{\circ}$ in Gaia.
@@
Figure 17: Visualized stream catalog in the sky, where color scale indicates the the minimum detectable subhalo mass, $M_{\mathrm{sh}}^{\mathrm{min}}$, with respect to observational scenario “DESI+Gaia” (top) and “Via+Gaia+LSST10” (bottom). The five streams with the best $M_{\mathrm{sh}}^{\mathrm{min}}$ under each observational scenario are labeled. Streams shown in light gray do not have an estimated $M_{\mathrm{sh}}^{\mathrm{min}}$: these have missing $M_{\rm stellar}$, are too short, or the observable number of stars is too low.
5.1 Stream ranking
@@
We use the catalog of cold stellar streams from Bonaca & Price-Whelan (2025), which contains information on stream width $\sigma_{\theta}$, stream length $l$, heliocentric distance to the stream $r_{h}$, and stellar mass $M_{\mathrm{stellar}}$. Note that we use $r_{h}$ as the analog of $r_{0}$ in the analytic model, as this determines the distance to the observer and thus the number of observable stars. Of the 131 total streams in the catalog, we only consider the 86 streams with a reported stellar mass. Streams with length $l<20^{\circ}$ are also filtered out based on the analysis in Sec. 4.4. We estimate the total number of stars per unit angle as $\lambda\approx 4N_{\rm stars}/3l$ (see discussion around Eq. 12), where $N_{\rm stars}$ is obtained from $M_{\mathrm{stellar}}$ using the distribution of stars in the IMF. We also exclude streams whose observable number of stars, $n_{\mathrm{obs}}$, is less than 1 star per $2^{\circ}$ for the Gaia magnitude limit of $G<20.7$, since our sensitivity estimates break down in that limit. While such streams might have sufficiently large $n_{\mathrm{obs}}$ for Via or LSST observations, they still generally have worse statistics and therefore are sensitive only to higher subhalo masses.
@@
We present visualizations of the stream catalog and our results in Figs. 17-20. First, Fig. 17 summarizes our main result, showing all of the streams on the sky with color scale indicating subhalo detectability. The 5 most promising streams with respect to each observational scenario are also indicated.
@@
Fig. 18 provides more detailed information for individual streams. The left panel shows $M_{\mathrm{sh}}^{\mathrm{min}}$ for all scenarios and streams, where the streams are ranked by the best case observational scenario. The right panel, adapted from Bonaca & Price-Whelan (2025), shows how those streams appear on the sky in the same order (left to right, and top to bottom) as in the left panel. Each subplot shows sky positions of stream members (black points) and stream orbits (blue lines) in individual stream reference frames. Further details on how the stream members and orbit are obtained can be found in Appendix A and B of Bonaca & Price-Whelan (2025). Generally, the streams with the best prospects appear thin, dense, and long on the sky. However, the right panel also makes it clear that streams must be evaluated on a case-by-case basis, as some streams are not well described by a simple stream orbit and other effects can play an important role.
@@
Some of the streams with promising prospects are likely to be highly impacted by baryonic structures, which will limit their use for subhalo detection. For instance, Palomar 5 has features induced by interactions with the bar (Pearson et al., 2017). This is a result of the particular orbit associated with Palomar 5, which is prograde and passes close to the Galactic bar. In general, streams on orbits far from the Galactic Center or on retrograde orbits are less likely to be affected by the Galactic bar and more promising for subhalo detection. In Tab. 5, we have included a column “Retrograde/Prograde” for the orbital direction of the streams, but more study is needed to determine the impact of the bar on individual streams. Orphan-Chenab appears to be similarly promising in our simple model, but it has features which have been attributed to a close encounter with the Large Magellanic Cloud (Erkal et al., 2019). The effects above would have to be accounted for in searching for subhalo signatures.
@@
Another example where our model cannot capture all effects comes from Jhelum. In terms of subhalo detection, it is ranked relatively highly here, partly due to its narrow width $\sigma_{\theta}=0.33^{\circ}$. However, there is also a much thicker component associated with Jhelum as well, which may be evidence of some more complicated dynamics (Bonaca et al., 2019b). Furthermore, while we have assumed a fairly uniform stellar density along the streams, Jhelum illustrates how observed streams can exhibit many more fluctuations due to observational selection criteria in addition to real density variations.
@@
Fig. 19 illustrates the correlations in stream properties and prospects for subhalo detectability, for the best-case observational scenario. All three properties ($\lambda$, $\sigma_{\theta}$, $r_{0}$) play a role in determining detectability. The five streams with the best prospects are labeled. While high stellar density $\lambda$ and low width $\sigma_{\theta}$ are clearly important, there is no single stream property that can be used to assess detectability.

### v3

@@ 4.3 Confidence intervals on subhalo mass
In the denominator, $\hat{\bm{\theta}}$ refers to the values of the nuisance parameters which maximize the likelihood for that particular value of $M_{\rm sh}$. By definition, $t(M_{\mathrm{sh}})=0$ at the maximum likelihood point. In the large sample limit, the maximum likelihood point will be at $M_{\mathrm{sh}}^{\mathrm{true}}$ and the $n\sigma$ confidence interval on $M_{\mathrm{sh}}^{\mathrm{true}}$ is obtained by solving for the $M_{\rm sh}$ at which $t(M_{\mathrm{sh}})=n^{2}$.
@@
Figure 11: Test statistic $t(M_{\mathrm{sh}})$ used to obtain confidence intervals, as defined in Eq. 28. Here $t(M_{\mathrm{sh}})$ is evaluated on the Asimov data set for a stream with $\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$, $\lambda=100~\mathrm{deg}^{-1}$ and all the four observational scenarios. The dataset is generated from $M_{\mathrm{sh}}^{\mathrm{true}}=10^{8}\mathrm{M}_{\odot}$.
@@
[…]
@@ 4.4 Minimal stream length in angle
In addition to the stream properties used in Eq. 27, the stream length in angle, denoted by $l$, also affects detectability. So far, the default angular region in the likelihood calculations is $22^{\circ},22^{\circ}]$ around the impact location, for a total range of $\Delta\psi=44^{\circ}$ on a single stream arm. This region was selected to be able to capture the full perturbation profile of a fairly massive subhalo as high as $\sim 10^{9}~\mathrm{M}_{\odot}$. However, this assumes a stream which is at least twice the angular length, which would impose a strong requirement when considering Milky Way streams.
On the other hand, for the minimum detectable subhalo mass $M_{\rm sh}^{\rm min}$, the requirement on the stream length can be reduced significantly. We now consider variable angular regions of size $\Delta\psi/2,\Delta\psi/2]$ with total range $\Delta\psi$. In Fig. 13, we plot the median $M_{\rm sh}^{\rm min}$ as a function of $\Delta\psi$ for varying stream properties $\sigma_{\theta}$, $r_{0}$, and $\lambda$. For simplicity, we only consider the best observational scenario “Via + Gaia + LSST10” here. We follow the same procedure as that used to obtain Fig. 8. We only show a data point if the resulting $M_{\rm sh}^{\rm min}$ is smaller than $10^{8}~\mathrm{M}_{\odot}$.
@@ 4.5 Dependence on impact properties
  • Impact parameter $b$ Our default value of $b=0$ is the optimal case. The results for non-zero $b$ are shown in the top left panel of Fig. 14. The detectability remains almost the same when the impact parameter $b$ is below 100 pc, and then begins to degrade when $b$ goes beyond 100 pc. The turning point is at the same scale as the scale radius of a CDM subhalo of $10^{6}~\mathrm{M}_{\odot}$, which is 162 pc according to Eq. 25. This is consistent with our statement that the results do not vary significantly for impact parameter $b\lesssim r_{s}$, which can also be seen from the dependence of the velocity kicks on $b^{2}+r_{s}^{2}$ in Eqs. 1-3. Our results are limited to $b\leq 10^{3}$ pc given the assumptions of the analytic approximation, Eq. 4.
@@
    Figure 15: Decomposition of $q_{0}$ contributions from different observables. All settings are the same as the top panel of Fig. 9 except that we consider special times since flyby, $t=279.3$ Myr (top) and $t=558.6$ Myr (bottom), when the perturbation in the $z$ direction is 0. The contribution from $\overline{\Delta\theta}$ vanishes completely in these cases. Subhalo masses are chosen to be the minimum detectable subhalo mass, and $q_{0}$ is evaluated on the Asimov data set.
    As $t$ increases, the variations with $t$ become smaller. This is because the size of the gap caused by a subhalo impact grows with time, which makes the density observable $\rho/\rho_{0}$ more important over other periodic observables for detectability. The gap growth also causes a slight decrease in $M_{\rm sh}^{\rm min}$ as larger gaps are easier to detect. Fig. 15 also shows how the density observable plays a more important role as $t$ grows. Note that we do not consider much older impacts, which may be more difficult to detect. Impacts older than half the stream age tend to be located in the outer parts of the stream, where the stellar density is lower, and the perturbation may also be partially washed out by the stream dispersion. These effects are not captured in the analytic model.
  • Direction of subhalo velocity $\hat{w}$ In studying the effects of changing the absolute subhalo velocity $\bm{w}=(w_{r},w_{t},w_{z})$, we first keep its magnitude fixed at $w=180$ km/s, and look into the effect of different directions $\hat{w}$. Taking symmetry considerations into account, we pick the following subhalo directions as representatives: $\hat{z}~\mathrm{(default)},\hat{r},\pm\hat{t}+0.1\hat{z},\hat{z}+\hat{r},\hat{z}\pm\hat{t}$ and $\hat{r}\pm\hat{t}$, where $\hat{t}$ is direction along the stream motion. Note that velocities purely along $\pm\hat{t}$ break the assumption of the analytic model, Eq. 4, so we have added a small perpendicular component $0.1\hat{z}$ to them. The direction $\pm\hat{t}+0.1\hat{z}$ is consistent with the impulse approximation, with $\frac{w_{\mathrm{rel}}}{w_{\perp}}\sqrt{b^{2}+r_{s}^{2}}\lesssim 0.3r_{0}~~\text{and}~~\frac{V_{c}}{w_{\perp}}\sqrt{b^{2}+r_{s}^{2}}\lesssim 0.3r_{0}$ at the minimum detectable subhalo mass. The results are shown in the top right panel of Fig. 14.
@@ 5 Application to known streams
Name $\sigma_{\theta}$ $l$ $r_{h}$ $M_{\mathrm{stellar}}$ $\lambda$ Retro $_{\mathrm{sh}}^{\mathrm{min}} [M\mathrm{M}_{\odot}$] $_{\mathrm{sh}}^{\mathrm{min}} [M\mathrm{M}_{\odot}$] $_{\mathrm{sh}}^{\mathrm{min}} [M\mathrm{M}_{\odot}$] $_{\mathrm{sh}}^{\mathrm{min}} [M\mathrm{M}_{\odot}$]
[] [] [kpc] $\mathrm{M}_{\odot}$] [deg-1] /Prograde Gaia DESI + Gaia Via + LSST Via + LSST10
C-12 0.51 28 11.5 14000 528 P 6.75e+06 6.58e+06 1.40e+06 7.75e+05
ATLAS-Aliqa Uma 0.26 41 21.4 19000 490 P 1.05e+07 1.05e+07 1.65e+06 7.94e+05
300S 0.34 25 15.9 7600 321 R 1.16e+07 1.16e+07 2.10e+06 1.08e+06
NGC 6397 0.79 32 2.5 2500 83 P 2.85e+06 2.84e+06 1.38e+06 1.16e+06
Palomar 5 0.54 32 21.3 17000 561 P 2.23e+07 2.15e+07 3.20e+06 1.45e+06
GD-1 0.43 119 8.0 14000 124 R 9.06e+06 9.11e+06 2.39e+06 1.47e+06
Orphan-Chenab 1.02 137 20.7 130000 1003 P 2.79e+07 2.62e+07 3.77e+06 1.64e+06
Ylgr 0.72 49 9.5 11000 237 R 1.36e+07 1.33e+07 3.05e+06 1.72e+06
Gaia-6 0.4 21 8.3 1800 91 R 1.16e+07 1.17e+07 3.02e+06 1.84e+06
Kshir 0.23 37 10.7 2200 63 R 1.31e+07 1.36e+07 3.16e+06 1.86e+06
C-7 0.42 34 5.8 1500 47 R 1.07e+07 1.09e+07 3.43e+06 2.31e+06
NGC 5466 0.23 23 17.4 1900 87 R 2.53e+07 2.60e+07 4.59e+06 2.31e+06
Gaia-1 0.34 40 5.0 1100 29 R 9.22e+06 9.59e+06 3.32e+06 2.37e+06
Jhelum 0.65 97 13.0 17000 185 P 2.69e+07 2.65e+07 5.09e+06 2.60e+06
Leiptr 0.48 73 7.1 3000 43 R 1.95e+07 1.99e+07 5.49e+06 3.41e+06
Gaia-8 0.83 47 7.1 3600 81 R 2.24e+07 2.23e+07 5.90e+06 3.52e+06
Gaia-12 0.33 29 11.4 1200 44 R 3.07e+07 3.16e+07 6.85e+06 3.78e+06
SGP-S 0.21 29 9.5 530 19 P 2.48e+07 2.62e+07 6.52e+06 3.92e+06
New-25 0.88 23 50.3 18000 827 P 1.49e+08 1.40e+08 1.22e+07 4.02e+06
C-19 0.35 47 18.0 3200 72 P 5.21e+07 5.30e+07 8.83e+06 4.22e+06
Hrid 1.0 77 3.2 2000 27 R 1.49e+07 1.50e+07 6.14e+06 4.57e+06
Kwando 6.58 57 7.7 40000 742 P 4.96e+07 4.45e+07 9.62e+06 4.81e+06
Omega Centauri 4.25 81 5.2 20400 266 R 3.24e+07 3.00e+07 8.37e+06 4.86e+06
C-13 0.49 23 8.4 750 34 P 3.33e+07 3.40e+07 8.49e+06 4.96e+06
Sylgr 0.99 30 3.3 700 25 P 1.70e+07 1.72e+07 6.90e+06 5.08e+06
Phlegethon 0.93 77 3.4 1700 23 R 1.75e+07 1.78e+07 7.04e+06 5.16e+06
Indus 1.98 90 16.6 34000 399 P 8.57e+07 8.03e+07 1.23e+07 5.29e+06
C-24 0.58 37 14.8 2700 77 P 6.18e+07 6.19e+07 1.11e+07 5.38e+06
C-23 0.33 20 8.7 310 16 P 4.10e+07 4.29e+07 1.08e+07 6.37e+06
New-16 0.34 22 5.0 170 8 P 2.61e+07 2.77e+07 9.49e+06 6.61e+06
NGC 3201 0.75 111 4.9 2100 20 R 3.09e+07 3.16e+07 1.03e+07 6.82e+06
New-1 0.2 31 15.5 420 14 P 7.60e+07 8.09e+07 1.52e+07 7.77e+06
Slidr 1.67 35 3.1 840 25 P 2.75e+07 2.73e+07 1.09e+07 7.78e+06
NGC 288 1.23 30 11.0 2200 78 R 8.53e+07 8.36e+07 1.66e+07 8.30e+06
Gaia-9 1.49 36 4.2 950 28 R 3.96e+07 3.94e+07 1.33e+07 8.74e+06
M5 0.33 39 14.2 710 19 P 9.15e+07 9.55e+07 1.81e+07 9.16e+06
Gaia-11 0.46 50 12.5 1100 23 R 9.12e+07 9.39e+07 1.87e+07 9.57e+06
C-11 1.48 33 6.9 1100 35 P 8.36e+07 8.27e+07 2.10e+07 1.18e+07
New-7 0.76 26 2.8 110 4 P 3.68e+07 3.86e+07 1.71e+07 1.32e+07
C-9 1.71 25 5.4 630 27 P 7.83e+07 7.76e+07 2.24e+07 1.34e+07
M3 0.97 76 9.8 2000 28 P 1.20e+08 1.20e+08 2.59e+07 1.34e+07
New-10 1.0 26 2.3 100 4 P 3.78e+07 3.95e+07 1.91e+07 1.53e+07
M68 1.37 105 5.4 1700 17 P 8.65e+07 8.69e+07 2.55e+07 1.54e+07
New-13 0.75 29 6.2 250 9 R 9.23e+07 9.55e+07 2.71e+07 1.63e+07
C-25 2.13 25 11.0 1300 55 P 2.17e+08 2.10e+08 3.99e+07 1.87e+07
New-23 0.69 30 4.2 120 4 R 7.51e+07 7.91e+07 2.80e+07 1.91e+07
Wukong 5.99 63 18.1 15000 252 P 5.49e+08 5.01e+08 6.62e+07 2.48e+07
New-3 2.17 71 1.7 120 2 P 1.05e+08 1.08e+08 5.82e+07 4.70e+07
New-6 2.54 80 3.1 340 4 P 1.87e+08 1.89e+08 7.15e+07 4.76e+07
New-2 3.47 61 2.2 73 1 P 3.91e+08 4.00e+08 1.78e+08 1.25e+08
@@
Table 5: Milky Way streams and estimated minimum detectable subhalo mass at 95% CL under different observational scenarios. Stream properties are obtained from the catalog in Bonaca and Price-Whelan (2025). We only include streams with reported stellar mass, angular length $l>20^{\circ}$, and density of observable stars greater than 1 star per $2^{\circ}$ in Gaia.
@@
Figure 17: Visualized stream catalog in the sky, where color scale indicates the minimum detectable subhalo mass, $M_{\mathrm{sh}}^{\mathrm{min}}$, with respect to observational scenario “DESI+Gaia” (top) and “Via+Gaia+LSST10” (bottom). The six streams with the best $M_{\mathrm{sh}}^{\mathrm{min}}$ under each observational scenario are labeled. Streams shown in light gray do not have an estimated $M_{\mathrm{sh}}^{\mathrm{min}}$: these have missing $M_{\rm stellar}$, are too short, or the observable number of stars is too low.
@@
We use the catalog of cold stellar streams [^5] https://github.com/abonaca/streams_overview from Bonaca and Price-Whelan (2025), which is based on streams currently in the galstreams package (Mateu, 2023). The catalog contains information on stream width $\sigma_{\theta}$, stream length $l$, heliocentric distance to the stream $r_{h}$, and stellar mass $M_{\mathrm{stellar}}$. Note that we use $r_{h}$ as the analog of $r_{0}$ in the analytic model, as this determines the distance to the observer and thus the number of observable stars. Of the 131 total streams in the catalog, we only consider the 86 streams with a reported stellar mass. Streams with length $l<20^{\circ}$ are also filtered out based on the analysis in Sec. 4.4. We estimate the total number of stars per unit angle as $\lambda\approx 4N_{\rm stars}/3l$ (see discussion around Eq. 12), where $N_{\rm stars}$ is obtained from $M_{\mathrm{stellar}}$ using the distribution of stars in the IMF. We also exclude streams whose observable number of stars, $n_{\mathrm{obs}}$, is less than 1 star per $2^{\circ}$ for the Gaia magnitude limit of $G<20.7$, since our sensitivity estimates break down in that limit. While such streams might have sufficiently large $n_{\mathrm{obs}}$ for Via or LSST observations, they still generally have worse statistics and therefore are sensitive only to higher subhalo masses.
@@
We present visualizations of the stream catalog and our results in Figs. 17-21. First, Fig. 17 summarizes our main result, showing all of the streams on the sky with color scale indicating subhalo detectability. The 6 most promising streams with respect to each observational scenario are also indicated.
@@
Fig. 18 shows $M_{\mathrm{sh}}^{\mathrm{min}}$ for all scenarios and streams, where the streams are ranked by the best case observational scenario. Fig. 19, adapted from Bonaca and Price-Whelan (2025), shows how those streams appear on the sky in the same order (left to right, and top to bottom) as Fig. 18. Each subplot shows sky positions of stream members (black points) and stream orbits (blue lines) in individual stream reference frames. Further details on how the stream members and orbit are obtained can be found in Appendix A and B of Bonaca and Price-Whelan (2025). Generally, the streams with the best prospects appear thin, dense, and long on the sky. However, the sky plots also make it clear that streams must be evaluated on a case-by-case basis, as some streams are not well described by a simple stream orbit and other effects can play an important role.
Figure 18: Visualized stream catalog ranked by detectability under Via + Gaia + LSST10.
@@
Some of the streams with promising prospects are likely to be highly impacted by baryonic structures, which will limit their use for subhalo detection. For instance, Palomar 5 has features induced by interactions with the bar (Pearson et al., 2017; Erkal et al., 2017; Banik and Bovy, 2019). This is a result of the particular orbit associated with Palomar 5, which is prograde and passes close to the Galactic bar. In general, streams on orbits far from the Galactic Center or on retrograde orbits are less likely to be affected by the Galactic bar and more promising for subhalo detection. In Tab. 5, we have included a column “Retrograde/Prograde” for the orbital direction of the streams, but more study is needed to determine the impact of the bar on individual streams. Orphan-Chenab appears to be similarly promising in our simple model, but it has features which have been attributed to a close encounter with the Large Magellanic Cloud (Erkal et al., 2019). The effects above would have to be accounted for in searching for subhalo signatures.
@@
Another example where our model cannot capture all effects comes from Jhelum. In terms of subhalo detection, it is ranked relatively highly here, partly due to its narrow width $\sigma_{\theta}=0.65^{\circ}$. However, there is also a much thicker component associated with Jhelum as well, which may be evidence of some more complicated dynamics (Bonaca et al., 2019a). Furthermore, while we have assumed a fairly uniform stellar density along the streams, Jhelum illustrates how observed streams can exhibit many more fluctuations due to observational selection criteria in addition to real density variations.
@@
Fig. 20 illustrates the correlations in stream properties and prospects for subhalo detectability, for the best-case observational scenario. All three properties ($\lambda$, $\sigma_{\theta}$, $r_{0}$) play a role in determining detectability. The six streams with the best prospects are labeled. While high stellar density $\lambda$ and low width $\sigma_{\theta}$ are clearly important, there is no single stream property that can be used to assess detectability.